OpenStudy (annie12m):

Ill fan plus medal! Rewrite the expression by using grouping symbols. y • 4 + y • z A. y (4 + z) B. 4(y + z) C. z(y • 4) D. y • (4 + y) • z

RhondaSommer (rhondasommer):

which one gets multiplied on both sides?

OpenStudy (annie12m):

d?

RhondaSommer (rhondasommer):

no; i mean the equation

RhondaSommer (rhondasommer):

not for the answer

OpenStudy (annie12m):

y

RhondaSommer (rhondasommer):

so that means that y will be outside the parenthesis because of distirbutive property

OpenStudy (annie12m):

so the answer would be a

RhondaSommer (rhondasommer):

yup :)

OpenStudy (annie12m):

ok thanks, ill medal and fan

RhondaSommer (rhondasommer):

no prob :) tag me if you have anymore.

OpenStudy (annie12m):

will do

OpenStudy (annie12m):

can you check my work?? 7 (5 + 9) = 7 * x + 7 * 9 35 + 63 = 7 * x + 7 * 9 35 + 9 = x + 63 -32 + 9 = x -23 = x

OpenStudy (annie12m):

@RhondaSommer

RhondaSommer (rhondasommer):

here.

RhondaSommer (rhondasommer):

ummm...thats not quite correct

RhondaSommer (rhondasommer):

35 + 63 = 7x + 63 5 + 9 = x + 63 (divide both sides by 7) 14 = x + 63 (simplify) -49 = x (subtract both sides by 63)

RhondaSommer (rhondasommer):

i started were you made the mistake